OpenStudy (anonymous):

Identify the x- and y-intercepts of f(x). Without graphing g(x), identify its x- and y-intercepts.

OpenStudy (anonymous):

\[f(x)=-\frac{ 3 }{ 2 }x+9 ; g(x)=\frac{ 2 }{ 3 }f(x)\]

Parth (parthkohli):

x-intercept: the value of x when y = 0. y-intercept: the value of y when x = 0.

OpenStudy (anonymous):

y=9